OpenStudy (anonymous):

Help please!! Solve ln(5x + 7) = 8. Round to the nearest thousandth.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just a sec

OpenStudy (solomonzelman):

here is a trick for you: \(\large\color{black}{ \ln(5x+7)=8 }\) \(\large\color{black}{ \ln(5x+7)=8 \ln(e) }\) \(\large\color{black}{ \ln(5x+7)=\ln(e^8) }\) \(\large\color{black}{ 5x+7=e^8 }\) and on....

OpenStudy (solomonzelman):

see what I did?

OpenStudy (anonymous):

Let 5x+7=a ln(a)=8 Find a.

OpenStudy (solomonzelman):

or you could have just applied: \(\large\color{black}{ \log_ab=c\rightarrow a^c=b }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \log_ab=c~~~~~~~~~~~~~~~~\rightarrow~~~a^c=b }\) \(\large\color{black}{ \log_e(5x+7)=8~~~\rightarrow~~~e^8=5x+7 }\)

OpenStudy (solomonzelman):

subtract 7 from both sides, and divide both sides by 5. that will be the exact solution.

OpenStudy (anonymous):

e^8=a a=2.079

OpenStudy (solomonzelman):

And afterwards, you can approximate it, as you are told to.

OpenStudy (solomonzelman):

if I lost you, tell me where please.

OpenStudy (anonymous):

5x+7=2.079 5x=-4.921 x~~-0.984

OpenStudy (anonymous):

Sorta lost me at log a b

OpenStudy (solomonzelman):

Do you understand this rule? \(\large\color{black}{ \log_ab=c~~~~~~\Longrightarrow~~~~~a^c=b }\)

OpenStudy (solomonzelman):

without any reasons why, do you see what is going on in there?

OpenStudy (solomonzelman):

I'll color it, so you can better see.

OpenStudy (anonymous):

log 5x^7 = 8?

OpenStudy (solomonzelman):

\(\Large\color{black}{ \log_{\color{red}{a}}\color{green}{b}=\color{blue}{c}~~~~~~\Longrightarrow~~~~~\color{red}{a}^\color{blue}{c}=\color{green}{b} }\)

OpenStudy (solomonzelman):

you had: \(\Large\color{black}{ \ln{}\color{green}{(5x+7)}=\color{blue}{8} }\) which is same as, \(\Large\color{black}{ \log_{\color{red}{e}}\color{green}{(5x+7)}=\color{blue}{8} }\)

OpenStudy (solomonzelman):

apply the rule, now, what do you get?

OpenStudy (solomonzelman):

if you feel you don't get something, please ask. or if you want time to process, then don't even have to reply. take your time.

OpenStudy (anonymous):

I'm sorry about this, I'm just really bad with algebra :/ Um, e^8 = 5x + 7... and then just subtract 7 and divide 5 like you said earlier... -7/5 + e^8 = x ?

OpenStudy (solomonzelman):

okay, lets do this one together: \(\large\color{black}{ e^8=5x+7 }\) subtracting 7 from both sides (this is to isolate the 5x) \(\large\color{black}{ e^8\color{red}{-7}=5x+7\color{red}{-7} }\) the \(\large\color{black}{ 7 }\)s on the rigth side cancel. \(\large\color{black}{ e^8-7=5x }\)

OpenStudy (solomonzelman):

then divide both sides by 5... (try it, even if you feel uncomfortable doing that)

OpenStudy (anonymous):

\[\frac{ 5 }{ e^8 }-\frac{ 7 }{ 5 }=x\]

OpenStudy (anonymous):

or the e^8 on top :P

OpenStudy (solomonzelman):

it is \(\large\color{black}{ e^8 }\) over \(\large\color{black}{5 }\). like you just said, or asked, not sure about the intonation in that phrase.

OpenStudy (anonymous):

-1.4 as a decimal for -7/5. I don't know if that would help any..

OpenStudy (solomonzelman):

My connection was murdered

OpenStudy (anonymous):

Hm?

OpenStudy (solomonzelman):

Here. http://www.wolframalpha.com/input/?i=%28e%5E8%2F5%29-%287%2F5%29 note, how I put in the question/search, and how wolfram correctly interprets the input for me. and there is a "Decimal Approximation" it is right after the "Input", right below.

OpenStudy (solomonzelman):

if you have any question, ask....